Cameron invests $10,000 in an account that pays 5.3% per year. How much will he have if it is compounded weekly for 15 years?


1
Expert's answer
2021-03-11T12:26:50-0500

A = total amount (to be calculated)

P = $10,000 (principal or amount of money deposited)

r = 5.3% per year (annual interest rate)

n = 52 (number of times compounded per year)

t = 15 (time in years)


 To find the amount we use the formula:


"A=P(1+{r\over n})^{n.t}"

After plugging the given information we have


"A= 10000(1+\dfrac{0.053}{52})^{52\times15}"

"A=10000(1+0.001019231)^{780}"

"A=10000(1.001019231)^{780}"


"A=22135.47"

So, after 15 years Cameron has $22135.47